• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name lindgren

    Log entry time 20:45:17 on May11,2010

    Entry number 322199

    keyword=Rebooting Right Dipole NMR IOC

    19:30 4 ua in Hall A
    Megan and others are using beam to test Compton polarimeter
    Left dipole is still down
    Hareboot27 right dipole after several tries it looks like its working
    Field is fluctuating but NMR is locked???
    Jack Segal just came in
    Going to check left dipole controls and Q2 in the trailer.
    Jack says "The scope shows the NMR on the Right Dipole is locked in. The field readback is all over the map. I think we have fried a NMR meter in some fashion. If someone knows of a quirk in the meter that would do this please call and let us know. The current readback is very stable. Perhaps it would be best to set the current and just run hoping that the current readback is as accurate as we need."